# Interpretable Visualizations of Data Spaces for Classification Problems

### v1

@@ Abstract
How do classification models “see” our data? Based on their success in delineating behaviors, there must be some lens through which it is easy to see the boundary between classes; however, our current set of visualization techniques makes this prospect difficult. In this work, we propose a hybrid supervised-unsupervised technique distinctly suited to visualizing the decision boundaries determined by classification problems. This method provides a human-interpretable map that can be analyzed qualitatively and quantitatively, which we demonstrate through visualizing and interpreting a decision boundary for chemical neurotoxicity. While we discuss this method in the context of chemistry-driven problems, its application can be generalized across subfields for “unboxing” the operations of machine-learning classification models.
I Introduction
@@
Over the last many years, machine learning (ML) has transformed the way computational materials science research is being done, for example accelerating materials discovery[1], improving the quality of molecular simulations[2, 3], and enabling reliable predictions of materials properties[4]. However, many traditional ML and deep learning methods function as “black boxes,” providing little insight into the underlying physics and chemistry of materials and materials processes. However, in chemical research, the true value of an ML model often goes beyond its ability to make accurate predictions; it also lies in its ability to reveal how the set of input parameters, often carefully chosen to capture the underlying chemistry, relate to the property of interest. Addressing this need for interpretability has driven the development of approaches specifically tailored to uncover structure-property relationships, a pursuit central to the field of Quantitative Structure-Property Relationship (QSPR) modeling.
@@
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks[5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e. information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (U-MAP) [7], and Isometric Mapping (ISOMAP)[8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS)[10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
So, how do we determine which characteristics of our input data most influence the decisions made by our model? Feature importance algorithms numerically determine the impact of input features in determining the output. Tree-based methods, such as random forests, are most commonly associated with feature importance analysis, as they provide built-in “feature importance” measures, which are typically computed as the total reduction of either the Gini impurity (for classification) or the squared error (for regression) of each of the features [12]. SHAP (SHapley Additive exPlanations), another popular post-training analysis scheme, draws from game theory and involves assigning each feature a value (i.e. SHAP value) representing its contribution to a model’s prediction by considering all possible combinations of feature subsets [13]. Alternatively, many have measured the importance of different features by measuring the performance of feature subsets, either through iteratively adding or removing features[14]. These techniques can excel because they can identify the most relevant descriptors from a large feature set, but they may not be particularly informative when the features are heavily correlated or correspond to abstract concepts. In addition, the amount of information these methods provide beyond a ranking of the features is limited, and they tend to scale poorly when used with high-dimensional data[15].
Figure 1: Relationship between variables and projectors in principal covariates classification. $\mathbf{Z}$ represents the evidence, which is a quantification of class likelihoods, and it can be a tensor in a multilabel, multiclass classification problem.
@@
Thus, while methods from both unsupervised learning and supervised learning address aspects of the QSPR problem, neither can fully do so on their own. However, there exists a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning to more efficiently reveal structure-property relationships. Some of these methods include principal components regression [16], partial least squares regression [17], cluster-wise regression [18], continuum regression [19], and principal covariates regression (PCovR) [20]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective optimization problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [21], economic forecasting [22], psychology [23], bioinformatics [24, 25, 26], and molecular crystallization [27]. Similar to PCA, PCovR can also be used to incorporate nonlinear relationships through similarity kernels[28], and can be used with regularization to create sparse solutions [21, 25]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [29], farthest point sampling, and CUR decomposition [30].
@@
Despite its versatility and wide applicability, PCovR, along with most other hybrid learning methods, is primarily tailored for regression tasks (i.e. where the output is continuous). However, many important materials problems require us to determine relationships between design principles and discrete properties, as are common in classification problems. For classification, however, the lack of a method that integrates dimensionality reduction and classification performance makes QSPR difficult to perform. We propose a new method for QSPR, which we call principal covariates classification (PCovC), that involves using PCovR as an intermediary for predicting class likelihoods. We demonstrate the power of this approach by interpreting decision boundaries in chemical neurotoxicity data using various toxicity assays.
II Results
II.1 Mathematical Basis for Principal Covariates Classification
@@
In the following, we provide a concise description of the underlying mathematics for Principal Covariates Classification (PCovC). We direct interested readers to Helfrecht et al. [28] for a more thorough discussion, including that of data scaling and incorporating non-linearity. Included in the appendices are discussions of other QSPR techniques, namely SHAP analysis[13], linear discriminant analysis[31], and saliency maps[32].
II.1.1 Notation
@@
In this work we assume that the training data is organized into two objects: a ”feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the output for each observation. The feature matrix is formatted such that its rows correspond to the individual samples and that its columns correspond to the features. In other words, $\mathbf{X}$ has shape $n_{\mathbf{samples}}\times n_{\mathbf{features}}$. Likewise, the target matrix contains the output labels for each observation, each of which can take an integer value between 00 and $n_{\mathbf{classes}}-1$. Thus, $\mathbf{Y}$ has shape $n_{\mathbf{{samples}}}\times n_{\mathbf{{labels}}}$. We represent the class labels as integers. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\mathbf{samples}}\times n_{\mathbf{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\mathbf{samples}}\times n_{\mathbf{classes}}\times n_{\mathbf{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points.
In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P_{AB}}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write:
$\mathbf{T}$ $=\mathbf{XP_{XT}}$ (1)
$\mathbf{Z}$ $=\mathbf{TP_{TZ}}=\mathbf{XP_{XT}P_{TZ}}$ (2)
The relationship between these objects is shown in Fig. 1.
II.1.2 Linear Classification
@@
In a classification problem, the goal is to assign an input to one of $n_{\mathbf{classes}}$ discrete classes. In other words, we aim to draw “decision boundaries” that divide the input space into $n_{\mathbf{classes}}$ regions, with each region corresponding to a certain class. In a linear classification problem, these boundaries are lines, or hyperplanes if considering a higher-dimensional space. There are two main approaches to linear classification: deterministic and probabilistic. Deterministic approaches, such as support vector machines, involve finding a discriminant function and use an activation function to map this function to a specific class label. Probabilistic approaches, such as logistic regression, estimate the probability distributions of the input and output and use these probabilities to make classifications.
More formally, linear classifiers make predictions by establishing a linear projection of the input data:
$\mathbf{Z}=\mathbf{XP_{XZ}}$ (3)
and then using activation function $\phi$ to map the continuous $\mathbf{Z}$ to the set of discrete class labels $\mathbf{Y}\in{0,1,2,...n_{\mathbf{classes}}-1}$.
$\mathbf{Y}=\phi(\mathbf{Z})$
In the case of binary classification, we can determine the class by using a step function to apply a threshold to the elements in $\mathbf{Z}$. For example, in a binary support vector machine, the decision is based on the sign of $\mathbf{Z}$:
$\mathbf{Y}[i]=\begin{cases}1&\mathbf{Z}[i]>0\\ 0&\mathbf{else}\end{cases}$ (4)
II.1.3 Principal Covariates Regression
@@
PCovR[20, 28] is a mapping technique that determines a low-dimensional matrix $\mathbf{T}\equiv\mathbf{XP_{XT}}$ that minimizes the function
$\ell=\underbrace{\alpha||\mathbf{X-XP_{XT}P_{XT}^{T}}||^{\mathbf{2}}}_{\text{loss in reconstructing X}}+(1-\alpha)\underbrace{||\mathbf{Y-XP_{XT}P_{TY}}||^{2}}_{\text{loss in predicting Y}}$ (5)
For linear and kernel ridge regression, our projector from $\mathbf{T}$ ($\mathbf{P_{TY}}$) is an analytically solvable function of $\mathbf{X}$ and $\mathbf{Y}$, and thus Eq. (5) is analytically solvable by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$.
$\mathbf{\tilde{K}}=\mathbf{\alpha XX^{T}}+(1-\alpha)\mathbf{\hat{Y}\hat{Y}^{T}}$ (6)
or, equivalently, a modified covariance matrix $\mathbf{\tilde{C}}$
$\mathbf{\tilde{C}}=\alpha\mathbf{X}^{T}\mathbf{X}+(1-\alpha)\left(\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{X}^{T}\mathbf{\hat{Y}\hat{Y}}^{T}\mathbf{X}\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\right)$ (7)
@@
where $\mathbf{\hat{Y}}=\mathbf{XP}_{XY}$ is the linear regression approximation of $\mathbf{Y}$, used to prevent overfitting. This procedure is easily adapted for methods based on data similarity, i.e., kernel models, as detailed in Helfrecht et al. [28].
II.1.4 Principal Covariates Classification
@@
The PCovR loss function (Eq. 5) should not be directly modified to be used in classification tasks – replacing the regression loss with a classification loss often results in a non-analytical function (when the underlying classification algorithm incorporates activation functions into their loss optimization) or is scientifically non-productive, as many classification tasks, particularly multiclass tasks, use numerical values to denote class, where these numerical values are ostensibly arbitrary. We choose to adapt PCovR for classification tasks by leveraging the evidence as our approximation of $\mathbf{Y}$. This yields a new modified Gram matrix $\mathbf{\tilde{K}}$.
@@
whose eigendecomposition (or that of the modified covariance matrix, analogous to Eq. 7) can be used to obtain a relevant low-dimensional map. Here, $\mathbf{Z}$ can be obtained through extracting the weights from a fitted linear classifier and applying Eq.  3. Similar to PCovR, this can be easily adapted for classification models operating on data similarity through kernel methods.
@@
II.2 Exposing decision boundaries in toxicity classifications
Figure 2: Effect of mixing parameter $\alpha$ on the low-dimensional map and classification performance (on an out-of-sample dataset) for the “tox21-ache-p3 (-1)” assay. White points indicate molecules deemed “non-toxic” (Negative) by the assay, blue “toxic” (Positive). Opacity denotes train / test split, with low-opacity points belonging to the training set. In the bottom right of each panel is a confusion matrix showing the accuracy of logistic regression on the resulting map, where “TN” indicates the number of “true negatives”, “FP” indicates the number of “false positives”, and so on. For comparison, logistic regression on the full-dimensional data obtains a confusion matrix of TP=17, TN=505, FP=19, FN=19.

### v2

@@ Abstract
How do classification models “see” our data? Based on their success in delineating behaviors, there must be some lens through which it is easy to see the boundary between classes; however, our current set of visualization techniques makes this prospect difficult. In this work, we propose a hybrid supervised-unsupervised technique distinctly suited to visualizing the decision boundaries determined by classification problems. This method provides a human-interpretable map that can be analyzed qualitatively and quantitatively, which we demonstrate through several established examples from literature. While we discuss this method in the context of chemistry-driven problems, its application can be generalized across subfields for “unboxing” the operations of machine-learning classification models.
1 Introduction
@@
Over the past decades, machine learning (ML) has transformed computational chemistry, for example accelerating materials discovery [1], improving the quality of molecular simulations [2, 3], and enabling reliable predictions of materials properties [4]. However, many traditional ML and deep learning methods function as “black boxes,” providing little insight into the underlying physics and chemistry of materials and materials processes. However, in chemical research, the true value of an ML model often goes beyond its ability to make accurate predictions; it also lies in its ability to reveal how the set of input parameters, often carefully chosen to capture the underlying chemistry, relate to the property of interest. Addressing this need for interpretability has driven the development of approaches specifically tailored to uncover structure-property relationships, a pursuit central to the field of Quantitative Structure-Property Relationship (QSPR) modeling.
@@
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks [5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e., information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (U-MAP) [7], and Isometric Mapping (ISOMAP) [8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS) [10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
@@
So, how do we determine which chemical or materials characteristics most influence the decisions made by our model? When these characteristics are explicitly included in our input vectors, feature importance algorithms are indispensable for such tasks. Tree-based methods, such as random forests, are most commonly associated with feature importance analysis, as they provide built-in “feature importance” measures, which are typically computed as the total reduction of either the Gini impurity (for classification) or the squared error (for regression) of each of the features [12]. SHAP (SHapley Additive exPlanations), another popular post-training analysis scheme, draws from game theory and involves assigning each feature a value (i.e., SHAP value) representing its contribution to a model’s prediction by considering all possible combinations of feature subsets [13]. Alternatively, many have measured the importance of different features by measuring the performance of feature subsets through iteratively adding or removing features [14]. These techniques can excel because they can identify the most relevant descriptors from a large feature set. Still, they may not be particularly informative when the features are heavily correlated or correspond to abstract concepts.
@@
Thus, while methods from unsupervised learning and supervised learning address aspects of the QSPR problem, none can fully do so on their own. To address this problem, there exists a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning. Some of these methods include principal components regression [16], partial least squares regression [17], cluster-wise regression [18], continuum regression [19], and principal covariates regression (PCovR) [20]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [21], economic forecasting [22], psychology [23], bioinformatics [24, 25, 26], and molecular crystallization [27]. Similar to PCA, PCovR can also be used to incorporate non-linear relationships through similarity kernels [28], and can be used with regularization to create sparse solutions [21, 25]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [29], farthest point sampling, and CUR decomposition [30].
@@
Despite its versatility and wide applicability, PCovR, along with most other hybrid learning methods, is primarily used for regression tasks, where the output is continuous. Yet, many important materials problems require us to determine relationships between design principles and discrete properties, as are common in classification problems. For classification, however, the lack of a method that integrates dimensionality reduction and classification performance makes QSPR difficult to perform. We propose a new method for QSPR, which we call principal covariates classification (PCovC), that involves using PCovR as an intermediary for predicting class likelihoods.
2 Theory
@@
We provide a concise description of the underlying mathematics for principal covariates classification (PCovC). We direct interested readers to [28] for a more thorough discussion, including that of data scaling and incorporating non-linearity. Included in App. B are discussions of other QSPR techniques, namely SHAP analysis [13], linear discriminant analysis [31], and saliency maps [32].
2.1 Notation
@@
We assume that the data is organized into two objects: a “feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the classes for each observation. The rows of the feature matrix $\mathbf{X}$ correspond to the individual samples and the columns correspond to the features, leading to shape $n_{\text{samples}}\times n_{\text{features}}$. The target matrix $\mathbf{Y}$ has shape $n_{\text{{samples}}}\times n_{\text{{labels}}}$ and contains the output labels for each observation, each of which can take an integer value between 00 and $n_{\text{classes}}-1$. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\text{samples}}\times n_{\text{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\text{samples}}\times n_{\text{classes}}\times n_{\text{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points. In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P}_{AB}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write $\mathbf{T}\equiv\mathbf{XP}_{XT}$ and $\mathbf{Z}=\mathbf{TP}_{TZ}=\mathbf{XP}_{XT}\mathbf{P}_{TZ}.$ The relationship between these objects is shown in Fig. 1.
Figure 1: Relationship between variables and projectors in principal covariates classification. $\mathbf{Z}$ is the evidence matrix, which is a quantification of class likelihoods, and it can be a tensor in the case of a multilabel, multiclass classification problem.
2.2 Linear Classification
@@
The goal of a classification model is to assign an unseen sample to one of $n_{\text{classes}}$ discrete classes. Thus, we aim to draw “decision boundaries” that divide the input space into $n_{\text{classes}}$ regions, with each region corresponding to a certain class. In a linear classification problem, these boundaries are lines or hyperplanes in a higher-dimensional space. Deterministic approaches, such as support vector machines, involve finding a discriminant function and use an activation function to map this function to a specific class label. Probabilistic approaches, such as logistic regression, estimate the probability distributions of the input and output and use these probabilities to make classifications. More generally, linear classifiers make predictions by establishing a linear projection of the input data $\mathbf{Z}=\mathbf{XP}_{XZ}.$ and then using some function $\phi$, typically called an activation function, to map the continuous $\mathbf{Z}$ to the set of discrete class labels $\mathbf{Y}\in{0,1,2,...n_{\text{classes}}-1}$, such that $\mathbf{Y}=\phi(\mathbf{Z}).$
Linear classifiers are not typically used for visualization, with one strong exception: linear discriminant analysis, or LDA. LDA determines a linear projector $\mathbf{P}_{XT}$ of a given dataset that optimizes the ratio of inter-cluster scatter $\mathbf{S}_{B}\equiv\sum_{c}N_{c}\left(\langle\mathbf{X}_{c}\rangle-\langle\mathbf{X}\rangle\right)\left(\langle\mathbf{X}_{c}\rangle-\langle\mathbf{X}\rangle\right)^{T}$ and intra-cluster scatter $\mathbf{S}_{W}\equiv\sum_{c}\sum_{i\in c}\left(\mathbf{x}_{i}-\langle\mathbf{X}_{c}\rangle\right)\left(\mathbf{x}_{i}-\langle\mathbf{X}\rangle_{c}\right)^{T}$, where $N_{c}$ is the number of points in cluster $c$, $\langle\mathbf{X}_{c}\rangle$ is the column-wise average of features of points belonging to cluster $c$, and $x_{i}$ is the feature vector for point $i$ belonging to cluster $c$. The optimization function $\mathcal{J}(\mathbf{P}_{XT})$ is then written as:
$\mathcal{J}(\mathbf{P}_{XT})=\frac{\left|\mathbf{P}_{XT}^{T}\mathbf{S}_{B}\mathbf{P}_{XT}\right|}{\left|\mathbf{P}_{XT}^{T}\mathbf{S}_{W}\mathbf{P}_{XT}\right|}$
This will yield an $n_{\text{classes}}-1$ projection of the data that can be analyzed for quantitative structure-property relations. We have included comparisons with LDA within the text to demonstrate the abilities and shortcomings of this method. Further discussion of LDA can be found in App. B.2.
2.3 Principal Covariates Regression
@@
PCovR [20, 28] is a mapping technique to determine a low-dimensional $\mathbf{T}\equiv\mathbf{XP}_{XT}$ that minimizes
$\ell=\alpha\underbrace{||\mathbf{X-XP}_{XT}\mathbf{P}_{XT}^{T}||^{2}}_{\text{loss in reconstructing X}}+(1-\alpha)\underbrace{||\mathbf{Y}-\mathbf{XP}_{XT}\mathbf{P}_{TY}||^{2}}_{\text{loss in predicting Y}},$ (1)
@@
where a mixing parameter $\alpha$ is used to determine the relative weight of data reconstruction versus property regression. For linear and kernel ridge regression, our projector from $\mathbf{T}$ ($\mathbf{P}_{TY}$) is an analytically solvable function of $\mathbf{X}$ and $\mathbf{Y}$, and thus Eq. (1) is analytically solvable by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}=\alpha\mathbf{XX}^{T}+(1-\alpha)\mathbf{\hat{Y}\hat{Y}}^{T},$ or, equivalently, a modified covariance matrix $\mathbf{\tilde{C}}=\alpha\mathbf{X}^{T}\mathbf{X}+(1-\alpha)\left(\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{X}^{T}\mathbf{\hat{Y}\hat{Y}}^{T}\mathbf{X}\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\right),$ where $\mathbf{\hat{Y}}=\mathbf{XP}_{XY}$ is the linear regression approximation of $\mathbf{Y}$, used to prevent overfitting. This procedure is easily adapted for methods based on data similarity, i.e., kernel models, as detailed in [28].
2.4 Principal Covariates Classification
@@
The PCovR loss function, Eq. (1), should not be directly modified to be used in classification tasks – replacing the regression loss with a classification loss often results in a non-analytical function (when the underlying classification algorithm incorporates activation functions into their loss optimization) or is scientifically non-productive, as many classification tasks, particularly multiclass tasks, use numerical values to denote class, where these numerical values are ostensibly arbitrary. We choose to adapt PCovR for classification tasks by leveraging the evidence as our approximation of $\mathbf{Y}$. This yields a new modified Gram matrix (or modified covariance matrix)
@@
whose eigendecomposition can be used to obtain a relevant low-dimensional map. Because different classification algorithms operate to optimize different loss terms, they will yield different values of $\mathbf{Z}$; this is visualized in Fig. 2, where we demonstrate the new latent space maps determined for PCovC at $\alpha=0.5$ for the archetypal Iris dataset [33, 34] and several popular linear classification algorithms.
Similar to PCovR, PCovC will inherit the regularization incorporated into its parent supervised algorithm, which will prevent overfitting, and can be trivially adapted for non-linear kernel classification models, where instead $\mathbf{Z}\equiv\mathbf{KP}_{KZ}$. As written, this approach is not well-suited for decision trees or ensemble classification methods, which require multiple analogs of the evidence matrix.
Figure 2: PCovC maps will change based on the underlying classifier. PCovC ($\alpha=0.5$) maps for the Iris dataset [33, 34] with ridge classification, logistic regression, support vector classification, and a single-layer perceptron. Color corresponds to the class (red, green, and blue for setosa, versicolor, and virginica flower classes) and the background shows the estimated decision boundaries.
3 Results and Discussion
In the following case studies, we aim to show both the utility and flexibility of this approach across many different topic areas, data representations, and classification architectures. While we aim to demonstrate the extraction of actionable and insightful principles from each of these studies, we encourage the readers to refer to the original articles for more thorough scientific discussion.
@@ 3 Results and Discussion
3.1 What makes a molecule toxic? Exposing decision boundaries in toxicity classifications